# WildSAT: Learning Satellite Image Representations from Wildlife Observations

### v1

@@ 4.1 Satellite Image Classification
Table 1: Results of linear probing different models on seven downstream datasets without (Base) and with (+WS) WildSAT fine-tuning. Accuracy is reported for all datasets except BEN20k that reports micro F1 score. Base refers to the original models specified as the encoder and +WS refers to the same models further trained on the species observation data. Fine-tuning models with species observation show significant improvement over the base models. Both Prithvi-100M and SatCLIP are pre-trained with multispectral images, but for consistency across downstream datasets and models, only RGB bands are used.
TaxaBind GRAFT CLIP WildSAT
[60] [45] [57] (Ours)
Average Performance 59.8% 65.0% 71.6% 76.6%
Table 2: Average linear probing performance across all seven satellite image classification datasets using models with CLIP ViT-B/16 as the base. TaxaBind and GRAFT fine-tune CLIP and use additional modalities (e.g. text, satellite images, ground images) for cross-modal tasks. We show that fine-tuning CLIP with WildSAT improves performance on both CLIP and other CLIP-based models. We include accuracy per dataset in the Appendix.
4.2 Bird Species Encounter Rate Prediction
We further evaluate our satellite image representations by predicting species encounter rates. SatBird [64] introduces a benchmark for predicting bird species encounter rates for an area given the satellite image. Encounter rates for a location are given as a vector $\mathbf{r}\in[0,1]^{S}$ for $S$ species, where the $i$-th element is the probability for a visitor to observe species $i$ at that location. Three subsets are defined based on the location and season: (1) USA summer, (2) USA winter, and (3) Kenya. For Kenya, all bird species are considered year-round since migration is negligible. Models are evaluated using top-$k$ accuracy, which takes the top $k$ predicted encounter rates in an area and compares the corresponding species with the ground truth species present in the area. $k$ is the number of actual species observed in an area.
4.3 Base Models
Base models refer to the different pre-trained encoders before we fine-tune with WildSAT. We experiment on seven pre-training methods spanning random initialization, in-domain pre-training, and out-of-domain pre-training. These cover different architectures ResNet50, Swin-T, Swin-B, ViT-B/16, and ViT-L/16 for a total of 18 base models.
ImageNet [13] models are pre-trained with supervision on the updated ImageNet1K V2 dataset [58]. Previous works use V1 of the dataset [13], but we opted for the updated version which improves performance on the ImageNet benchmark by 3-4%. We include results on the ImageNet1K V1 pre-trained models in the Appendix.
MoCov3 [6] models are pre-pretrained using self-supervision with InfoNCE [53] on ImageNet1K.
SeCo [46] models are self-supervised on time augmented Sentinel-2 satellite images.
Random denotes randomly initialized models.
5 Results and Discussion
5.1 Downstream Classification Performance
Table 1 displays the results on the 7 downstream classification datasets across 15 different models/architectures. The addition of WildSAT improves 98 out of the 105 settings with an overall average improvement ranging from 7.7% to 17.4% across the different datasets (4.3% to 10.4% without the randomly initialized models).
Figure 3: Zero-shot results for text-based satellite image retrieval. The columns show the top 5 images returned given the text query on top. A model can be queried using general landscape descriptions such as ‘desert’, ‘ocean’, ‘forest’, and ‘grassland’. In addition, specific wildlife text such as ‘ibex’ and ‘gull’ can be used as queries to view the types of environment they inhabit. The ‘ibex’ inhabits mountains and high elevation areas, ‘gulls’ are typically found near bodies of water, ‘american alligators’ are found in swamps and the coast, the ‘house finch’ is commonly found in urban areas, and ‘corals’ are in the ocean—consistent with the retrieved satellite images shown.
WildSAT improves satellite image representations. The results in Table 1 highlight the performance improvements WildSAT contributes. These improvements may be attributed to our use of diverse supervision—integrating images, text embeddings, and species data at scale. This strategy ultimately helps in downstream classification, particularly for both increasing true positive rates on classes related to habitats (e.g. forests, deserts), while reducing false positives on the same types of classes. We show how classes related to wildlife habitats improve performance in the Appendix. Meanwhile, Table 2 provides a comparison of WildSAT (CLIP ViT-B/16 as base) with TaxaBind [60] and GRAFT [45], both of which also fine-tune CLIP ViT-B/16 using cross-modal supervision from ground images. While the latter two methods show improvements on cross-modal tasks, their performance when linear probed on satellite imagery tasks suffers compared to CLIP, indicating a degree of “forgetting” of the representations. In contrast, our method not only outperforms standard CLIP but also achieves the best results overall.
Satellite pre-trained models see a larger boost in performance. The results demonstrate that training with WildSAT can improve performance by as much as 10% on satellite pre-trained models such as SeCo and SatlasNet. While models pre-trained on ImageNet and MoCov3 also see performance improvements, we see less improvements on the AID, RESISC45, and FMoW datasets. This could be attributed to the three datasets having less categories related to species habitats (e.g. storage tank, airport, church). Additionally, since ImageNet and MoCov3 pre-trained models are already performing effectively, there is little room for improvement. WildSAT is trained on a dataset that is specifically geared towards habitats and land characteristics. Thus, we see more improvements on So2Sat20k and BEN20k, which cover climate zones and land cover types.
Larger performance gains on ViTs than ResNets. Despite observing smaller improvements with the ResNet50 ImageNet and MoCov3 models, larger improvements are observed with their ViT counterparts. Consistent with observations from previous work [57], we see better performance with the addition of other modalities when using transformers. This could be attributed to the more flexible representations of ViTs, unlike CNNs that inherently incorporate a strong inductive bias from the use of filters. Using attention layers in ViTs likely makes their embeddings more adaptable to other modalities such as text and location.
Kenya USA Summer USA Winter
Base +WS Base +WS Base +WS
SatlasNet [4] 23.90 24.40 48.59 50.03 54.02 55.01
SatMAE [12] 23.66 23.75 46.09 48.69 52.40 53.86
TaxaBind [60] 23.83 24.20 48.31 49.87 53.76 55.00
Average 23.80 24.12 47.66 49.53 53.39 54.62
UCM [74] AID [72] RESISC45 [7] So2Sat20k [78]
24.3% 93.2% 25.2% 84.4% 25.5% 88.2% 5.9% 41.8% 48.6%
44.2% 95.0% 41.6% 85.1% 43.0% 89.3% 18.3% 43.4% 57.5%
60.0% 95.0% 48.7% 86.2% 48.2% 88.8% 25.2% 44.2% 62.0%
70.0% 95.4% 55.6% 86.2% 58.9% 89.7% 27.9% 45.0% 66.1%
79.9% 97.2% 68.2% 88.9% 74.7% 93.0% 42.3% 55.2% 74.9%
5.2 Bird Species Encounter Rate Prediction
Table 3 shows WildSAT complementing existing methods on bird species encounter rate prediction. We take models SatlasNet (Swin-B) [4] and SatMAE (ViT-L/16) [12] presented in SatBird [64] that use satellite images to predict encounter rates of all bird species in an area. TaxaBind [60] is also added as a baseline. Using WildSAT improves these existing methods by 1-2%.
5.3 Zero-shot Image Retrieval
When trained using our WildSAT framework, we observe that models learn wildlife-specific attributes. By using the frozen satellite image encoder and a large language model, a user can input text to query satellite images. The top $k$ images with the most similar embeddings to the text embeddings (computed using cosine similarity) can be retrieved (Figure 2b.2). Figure 3 displays examples of satellite images retrieved given different text queries. General descriptions of landscapes or locations can be used for querying such as ‘desert’, ‘ocean’, ‘forest’, or ‘grassland’. At the same time, specific wildlife text can also be used as queries such as ‘ibex’ or ‘gull’. We see that zero-shot retrieval returns images of the habitat of the corresponding wildlife. This feature could potentially be used to find habitats for species with limited observation data, and can serve as a reference for species distribution studies.
5.4 Ablations
Our WildSAT framework is composed of multiple components: a satellite image encoder, location encoder, and text encoder. We investigate the contribution of each of these components in an ablation study. In Table 4 we ablate two models: a randomly initialized ResNet50 and an ImageNet pre-trained ViT-B/16. Columns with check marks on ‘loc’ and/or ‘env’ indicate the use of the location encoder ($\mathcal{L}_{\text{loc}}$ term in Eqn. 1). The location encoder can either use only location as input or use both location and environmental covariates. The ‘text’ and ‘img-a’ columns indicate the use of the large language model ($\mathcal{L}_{\text{txt}}$) and contrastive loss on the augmented satellite images ($\mathcal{L}_{\text{img}}$), respectively. A complete ablation of all possible combinations of WildSAT components for other models is also provided in the Appendix.
Each modality enhances performance. Larger improvements are observed for the randomly initialized model, since it starts from a lower accuracy and there is a larger room for improvement. Nonetheless, similar improvement trends are seen with both types of models. Improvements from location and environmental covariates can be from associating satellite images with particular locations and corresponding environmental characteristics (i.e. satellite images are mapped to specific latitudes and longitudes that are along the coast, or in the mountains). The addition of text further improves the performance while enabling zero-shot image retrieval capabilities. Text provides a rich source of information with more detailed descriptions of areas. Adding the satellite image loss term likely improves general image understanding, such as the model learning the rotation invariance of satellite images. Overall, each component of WildSAT contributes to performance improvements.
Different modalities can strengthen models by covering model-specific gaps. We hypothesize that models trained on satellite imagery datasets benefit primarily from location and text supervision associated with wildlife observation. This hypothesis is supported by observations that self-supervised models such as SeCo, SatMAE, and Prithvi—trained with objectives similar to WildSAT’s image self-supervision term on satellite datasets—still achieve significant gains (Table 1). Similarly, models like SatlasNet, which are trained with large-scale supervised learning on satellite images, also benefit. On the other hand, ImageNet pre-trained models benefit from the additional satellite image supervision. These results highlight the complementary nature of WildSAT’s supervision compared to existing datasets, which primarily focus on anthropogenic labels.
5.5 Limitations
While we show that WildSAT improves satellite image representations and has promising zero-shot performance, the datasets we train on contain inherent limitations that could affect model use. The US and Europe are overrepresented in the data as most citizen contributed data currently come from these locations. Underrepresented areas in Asia, Africa, Australia, and South America are likely to see less accurate results especially on satellite image retrieval. Satellite images additionally do not cover the extreme north and south of the Earth including Antarctica, so species endemic in those areas and their habitats might not be well-represented. We also note that observations of endangered wildlife were excluded or significant noise was added to their locations to ensure their safety. LLMs are further prone to generating hallucinations, which could impact model output reliability.
6 Conclusion
Acknowledgements.
Supplementary Material
Appendix A Datasets
A.1 Training Data Distribution
Figure A1 shows the spatial distribution of all data we collected across the globe. Most of the data are from United States and Europe, corresponding to the wildlife observation data available from citizen science platforms [2, 1].
A.2 Satellite Image Classification Evaluation Datasets
Below we briefly describe the different satellite image classification datasets used for evaluation.
UCM [74] contains 21 classes with 100 each covering USA. Each image is 256$\times$256 with a resolution of 1 ft.
AID [72] contains 30 class, each containing from 220 to 400 images from Google Earth. Each image is 600$\times$600 with a resolution ranging from 0.5 to 8 m.
RESISC45 [7] contains 45 classes with 700 images each, sourced from Google Earth. An image is 256$\times$256 with a resolution ranging from 0.2 to 30 m.
FMoW [8] contains 63 classes from over 200 countries with a total of over 400k images from QuickBird, GeoEye, and WorldView satellites. Images vary in size and resolution and each class has a different number of images.
EuroSAT [27] contains 10 classes of land use and land cover from Europe. Each image is 64$\times$64 with 10 m resolution. The dataset has 27k images with each class having a different number of images.
So2Sat20k [78] contains 17 classes across different climate zones with global coverage. The full dataset contains 400k pairs of Sentinel-1 and Sentinel-2 images. We use the GEO-Bench [37] version referred to as “So2Sat20k” which contains 20k training samples.
Appendix B Additional Implementation Details
WildSAT Training Details.
We train each base model on the WildSAT framework for 25 epochs using an Adam optimizer with a learning rate of $1\times 10^{-4}$, with an embedding dimension $d=512$, and a batch size of 64. Each satellite image is paired with a wildlife observation location. For each of these pairs, a section of text is randomly sampled from the Wikipedia [3] page of the species. A satellite image of the same location, but from a different time, is also randomly sampled for image augmentation. Random cropping, resizing, jitter, and channel mixing are applied as augmentations.
Satellite Image Filtering.
All satellite images are from Sentinel-2A and Sentinel-2B. We follow the same data collection procedure from SatlasPretrain [4], where satellite images are downloaded from EU’s Sentinel Data [17]. Each image is 512$\times$512 pixels with a 10 m resolution per pixel. Only images that are tagged with significantly less cloud cover from [4] are used. In addition, we only use satellite images that were taken in the same time range as the wildlife observation data (from 2017 to 2021). This is done since we do not use the exact observation date and time as an input to the model; we consider all observations throughout the time range. At the same time, the text descriptions we use also refer to all types of habitats regardless of time of year.
Multi-spectral Baselines.
Prithvi-100M [31] and SatCLIP [34] originally use multi-spectral data in their pre-training. However, for general applicability and easy comparisons with other models, we only use RGB bands. When WildSAT is applied to these models, we only fine-tune with the three bands, and set other bands to zero. At the same time, when applying both the base models and WildSAT fine-tuned models on downstream satellite image datasets, we also set other bands to zero.
Appendix C Additional Results
@@
Table A2: Linear probing results on downstream satellite classification datasets using models with CLIP as the base. Results are reported as accuracy, except for BEN20k which uses micro F1. TaxaBind and GRAFT both fine-tune a CLIP backbone and use additional modalities such as text, ground images, and satellite images for cross-modal tasks. WildSAT outperforms both the standard CLIP and the previous methods that also fine-tune on CLIP.
@@ C.1 Satellite Image Classification
Table A1 shows results on the ImageNet V1 base model that previous works have used [46, 44]. The results in Table 1 in the main paper include a base model using ImageNet V2 (also included in Table A1 for reference) which has generally better performance across the downstream satellite image datasets. Table A1 additionally shows results on SatMAE [12], a ViT-L/16 model that was pre-trained with the MAE framework on satellite images. Similar to previous results, WildSAT improves performance across the seven satellite image classification datasets evaluated.
@@
Table A2 displays the result for each evaluation dataset across different CLIP-based models. All models in the table starts with a pre-trained CLIP ViT-B/16 model [57]. TaxaBind[60] and GRAFT [45] use additional modalities such as ground images, text, and audio to improve model performance on cross-modal tasks such as zero-shot image-text retrieval. However, we show that while these same models do well on zero-shot tasks, they tend to “forget” some of the original image representations, with linear probing performance on downstream datasets lower than that of the original CLIP model. With WildSAT applied to CLIP, we show that we can outperform not only other CLIP-based models, i.e. GRAFT and TaxaBind, but also outperform the standard CLIP model across all the datasets in the evaluation. We hypothesize we can prevent “forgetting” by applying parameter efficient fine-tuning on out-of-domain pre-trained models such as CLIP. We further show this in Table A5.
@@
Figure A2 shows the confusion matrix of a sample result on the So2Sat20k test set. The matrix on the left shows the result of a base ImageNet pre-trained model, and the right matrix shows the result when WildSAT is applied. We show that WildSAT improvements on the true positive counts along the diagonal are largely due to fewer false positives on habitat-related classes. Looking at the second row of both matrices (class ‘Sparsely built’), the true positive count doubled from 24 to 48. A lot of this improvement comes from less false positives on ‘Scattered trees’ (from 7 false positives to 0), ‘Bare rock or paves’ (from 3 false positives to 0), and ‘Dense trees’ (from 1 false positive to 0)—all of which are habitat-related attributes. Similar trends can be observed on other classes as well such as in ‘Dense Trees’ and ‘Low Plants’ where we see higher true positive counts with WildSAT.
@@ C.2 Additional Ablations
Table A3 displays an ablation study conducted on two different types of models: a ResNet50 SeCo [46] model and a ViT-B/16 ImageNet [13] model. SeCo is pre-trained with a contrastive objective on time augmented satellite images (i.e. satellite images from the same location, but from different seasons). This objective is similar to the $\mathcal{L}_{\text{img}}$ term (Eqn. 1 in the main paper) in the loss function of our WildSAT framework. Thus, we see from Table A3, that simply adding the image augmentation term (‘img-a’) only slightly improves the average performance across the downstream datasets (from 70.1% to 71.8%). This small improvement could be attributed to the additional examples related to habitats that are possibly not as well-represented in the SeCo dataset. However, if we add other modalities such as text and location (in addition to the satellite image augmentation), we see a larger improvement with an average performance of 74.2% and 75.2%, respectively. In contrast, an ImageNet pre-trained model benefits from satellite image augmentations ($\mathcal{L}_{\text{img}}$ or ‘img-a’) since it was trained on a different domain. Simply adding the image augmentation term improved average performance from 76.9% to 82.3%. Further adding other modalities such as text and location also pushes the performance higher to 83.6%. These results support our hypothesis that WildSAT can complement and further improve different architectures by leveraging the different modalities.
@@
Table A4 shows an ablation study conducted on our choice of the location encoder. We use a ResNet50 SeCo encoder as the base model, and report accuracy on downstream classification datasets. All rows in the table use WildSAT with satellite images and text ($\mathcal{L}_{\text{img}}+\mathcal{L}_{\text{txt}}$), which are matched based on the location—i.e. location is implicitly used in all the results, and we ablate which encoder to use for explicitly including location as an input. We replace the location encoder in our WildSAT framework with one of the following: no model (i.e. use the position encoded latitude and longitude and/or the raw environmental covariates vector), SatCLIP [34], or SINR [11]. We use the SatCLIP pre-trained location encoder that takes the latitude and longitude as an input. For SINR, we explore the two variants of using (1) just the location (‘loc’) or (2) both the location (‘loc’) and environmental covariates (‘env’). We find that the best average performance uses SINR with both the location and environmental covariates.
@@
Table A5 displays the effect of fine-tuning all parameters of a given base model compared to fine-tuning specific layers (i.e. applying PEFT). We compare the effect on out-of-domain pre-trained models (e.g. ImageNet [58], CLIP [57]), and in-domain pre-trained models (e.g. SatlasNet [4], SeCo [46]). We find that out-of-domain pre-trained models have better downstream performance by applying scale and shift fine-tuning [20, 39], or by applying DoRa [47]. By fine-tuning specific layers, the models retain some of the original representations learned from the pre-training (i.e. ImageNet or CLIP) so that performance does not deteriorate compared to the base models. This has a significant impact on large models such as ViT, since fine-tuning all weights alters many parameters and risks shifting them in suboptimal directions. On the other hand, while applying PEFT for in-domain pre-trained models SatlasNet [4] and SeCo [46] improves performance compared to the base model, we see better performance when directly fine-tuning all the layers. This may be because the model has already undergone pre-training on satellite images, making additional pre-training on similar data from WildSAT result in a non-disruptive shift.

### v2

@@ Appendix C Additional Results
Table A1: Results of linear probing different models on seven downstream datasets without (Base) and with (+WS) WildSAT fine-tuning. Accuracy is reported for all datasets except BEN20k that reports micro F1 score. Base refers to the original models specified as the encoder and +WS refers to the same models further trained on the species observation data. Fine-tuning models with species observation show significant improvement over the base models. Both Prithvi-100M and SatCLIP are pre-trained with multispectral images, but for consistency across downstream datasets and models, only RGB bands are used. We include results on multispectral images in Tab. A4.
@@
Table A3: Linear probing results on downstream satellite classification datasets using models with CLIP as the base. Results are reported as accuracy, except for BEN20k which uses micro F1. TaxaBind, GRAFT, and RemoteCLIP fine-tune a CLIP backbone and use additional modalities such as text, ground images, and satellite images for cross-modal tasks. WildSAT outperforms both the standard CLIP and the previous methods that also fine-tune on CLIP.
Encoder So2Sat20k BEN20k
[84] [65]
Base +WS Base +WS
ViT-B/16 Prithvi-100M [32] 28.7 50.1 34.4 53.9
ViT-B/16 SatCLIP [35] 48.8 48.8 33.4 36.1
ResNet50 SatCLIP [35] 45.8 46.3 46.7 54.3
Average 41.1 48.4 38.2 48.1
Table A4: Linear probing results on models using multispectral images. Accuracy is reported for So2Sat20k and micro F1 score for BEN20k. ‘Base’ refers to the original models specified as the encoder and ‘+WS’ refers to the same models further trained with WildSAT. We see improved base model performance with using more bands (compared to using only RGB), and show that the addition of WildSAT further improves performance even with multispectral models and datasets.
@@ C.1 Satellite Image Classification
Linear Probing results.
Tab. A1 shows the raw numbers from which Fig. 3 was generated.
@@
Tab. A2 shows results on the ImageNet V1 base model that previous works have used [48, 46]. The results in Table 1 in the main paper include a base model using ImageNet V2 (also included in Tab. A2 for reference) which has generally better performance across the downstream satellite image datasets. Tab. A2 additionally shows results on SatMAE [13], a ViT-L/16 model that was pre-trained with the MAE framework on satellite images. Similar to previous results, WildSAT improves performance across the seven satellite image classification datasets evaluated.
Multi-spectral results.
Tab. A4 shows additional results when using multispectral images as input to the models that support this capability. Prithvi-100M and SatCLIP accept multiple bands as input, while So2Sat20k and BEN20k are datasets on GEO-bench that contain multiple bands. We find that WildSAT also improves on these multispectral models, demonstrating the broad applicability of our method.
@@
Tab. A3 displays the result for each evaluation dataset across different CLIP-based models. All models in the table starts with a pre-trained CLIP ViT-B/16 model [59]. TaxaBind[63] and GRAFT [47] use additional modalities such as ground images, text, and audio to improve model performance on cross-modal tasks such as zero-shot image-text retrieval. However, we show that while these same models do well on zero-shot tasks, they tend to “forget” some of the original image representations, with linear probing performance on downstream datasets lower than that of the original CLIP model. With WildSAT applied to CLIP, we show that we can outperform not only other CLIP-based models, i.e. GRAFT and TaxaBind, but also outperform the standard CLIP model across all the datasets in the evaluation. We hypothesize we can prevent “forgetting” by applying parameter efficient fine-tuning on out-of-domain pre-trained models such as CLIP. We further show this in Tab. A9.
@@
EuroSAT [28] So2Sat20k [84]
AnnualCrop SeaLake Highway Water Low plants Heavy industry
TaxaBind [63] 1.3 20.4 3.7 0.0 0.0 0.0
GRAFT [47] 33.3 72.1 55.6 16.4 0.0 13.4
RemoteCLIP [41] 19.6 15.9 5.3 36.5 5.9 0.0
WildSAT (Ours) 46.5 87.4 0.4 60.9 10.4 0.0
Table A5: Zero-shot image classification F1 score on different classes of EuroSAT and So2Sat20k with CLIP-based models. Since WildSAT is geared towards habitat-related classes, the coverage of zero-shot classification is less effective beyond natural concepts. WildSAT does well on natural classes like ‘SeaLake’ and ‘Water’, but struggles on anthropogenic classes like ‘Highway’ and ‘Heavy industry’.
Cashew1k [79] SAcrop3k [3]
Accuracy IoU Accuracy IoU
Base +WS Base +WS Base +WS Base +WS
ImageNet [14] 91.6% 91.9% 70.3% 70.6% 60.7% 62.3% 24.3% 25.0%
MoCov3 [7] 92.4% 93.2% 71.4% 73.3% 60.7% 60.8% 22.9% 24.9%
SeCo [48] 86.7% 93.2% 62.6% 73.3% 59.2% 59.4% 22.3% 22.8%
SatlasNet [5] 82.5% 91.8% 55.2% 71.0% 56.1% 57.1% 19.4% 20.5%
Random 69.8% 92.9% 40.1% 72.6% 54.7% 56.3% 18.0% 20.3%
84.6% 92.6% 59.9% 72.2% 58.3% 59.2% 21.4% 22.7%
Table A6: Downstream satellite image segmentation results. WildSAT can also improve on satellite segmentation tasks across different models and datasets. All models use the frozen pre-trained encoders with a convolutional-based decoder trained for each of the downstream tasks. ‘Base’ refers to the original models specified as the encoder and ‘+WS’ refers to the same models further trained on the species observation data.
Figure A3: Visualization of segmentation results using SatlasNet without and with WildSAT. WildSAT can more accurately identify classes ‘stable cashew plantations’ and ‘stable cropland/others’ in Cashew1k [79], and improve on the identification of ‘Rye’ and ‘Meadows’ in SAcrop3k [3]. Some areas of improvement are highlighted in red boxes.
@@
Fig. A2 shows the confusion matrix of a sample result on the So2Sat20k test set. The matrix on the left shows the result of a base ImageNet pre-trained model, and the right matrix shows the result when WildSAT is applied. We show that WildSAT improvements on the true positive counts along the diagonal are largely due to fewer false positives on habitat-related classes. Looking at the second row of both matrices (class ‘Sparsely built’), the true positive count doubled from 24 to 48. A lot of this improvement comes from less false positives on ‘Scattered trees’ (from 7 false positives to 0), ‘Bare rock or paves’ (from 3 false positives to 0), and ‘Dense trees’ (from 1 false positive to 0)—all of which are habitat-related attributes. Similar trends can be observed on other classes as well such as in ‘Dense Trees’ and ‘Low Plants’ where we see higher true positive counts with WildSAT.
WildSAT is effective on zero-shot classification of habitat-related classes.
Tab. A5 shows zero-shot image classification results on classes from EuroSAT and So2Sat20k. WildSAT is geared towards classes related to nature and animal habitats, and thus see improvements in those classes (e.g. ‘AnnualCrop’, ‘SeaLake’, ‘Water’, ‘Low plants’). While WildSAT underperforms on anthropogenic classes (e.g. ‘Highway’, ‘Heavy industry’) compared to models like TaxaBind and GRAFT, future work can integrate wildlife data and anthropogenic labels to create a more comprehensive and balanced representations.
Satellite image segmentation results.
Tab. A6 shows results of applying different ResNet50 encoders on satellite image segmentation. For each encoder, a convolution-based decoder is added while adopting the U-Net architecture [61]. To accurately evaluate the features learned with and without WildSAT, we freeze the encoder and fine-tune only the decoder on the downstream dataset. Tab. A6 shows that WildSAT can provide rich representations leading to better segmentation performance. Fig. A3 visualizes how WildSAT improves the segmentation outputs on Cashew1k and SAcrop3k. Further training on WildSAT leads to more accurate differentiation between the crop types.
@@ C.2 Additional Ablations
Tab. A7 displays an ablation study conducted on two different types of models: a ResNet50 SeCo [48] model and a ViT-B/16 ImageNet [14] model. SeCo is pre-trained with a contrastive objective on time augmented satellite images (i.e. satellite images from the same location, but from different seasons). This objective is similar to the $\mathcal{L}_{\text{img}}$ term (Eqn. 1 in the main paper) in the loss function of our WildSAT framework. Thus, we see from Tab. A7, that simply adding the image augmentation term (‘img-a’) only slightly improves the average performance across the downstream datasets (from 70.1% to 71.8%). This small improvement could be attributed to the additional examples related to habitats that are possibly not as well-represented in the SeCo dataset. However, if we add other modalities such as text and location (in addition to the satellite image augmentation), we see a larger improvement with an average performance of 74.2% and 75.2%, respectively. In contrast, an ImageNet pre-trained model benefits from satellite image augmentations ($\mathcal{L}_{\text{img}}$ or ‘img-a’) since it was trained on a different domain. Simply adding the image augmentation term improved average performance from 76.9% to 82.3%. Further adding other modalities such as text and location also pushes the performance higher to 83.6%. These results support our hypothesis that WildSAT can complement and further improve different architectures by leveraging the different modalities.
@@
Tab. A8 shows an ablation study conducted on our choice of the location encoder. We use a ResNet50 SeCo encoder as the base model, and report accuracy on downstream classification datasets. All rows in the table use WildSAT with satellite images and text ($\mathcal{L}_{\text{img}}+\mathcal{L}_{\text{txt}}$), which are matched based on the location—i.e. location is implicitly used in all the results, and we ablate which encoder to use for explicitly including location as an input. We replace the location encoder in our WildSAT framework with one of the following: no model (i.e. use the position encoded latitude and longitude and/or the raw environmental covariates vector), SatCLIP [35], or SINR [12]. We use the SatCLIP pre-trained location encoder that takes the latitude and longitude as an input. For SINR, we explore the two variants of using (1) just the location (‘loc’) or (2) both the location (‘loc’) and environmental covariates (‘env’). We find that the best average performance uses SINR with both the location and environmental covariates.
@@
Tab. A9 displays the effect of fine-tuning all parameters of a given base model compared to fine-tuning specific layers (i.e. applying PEFT). We compare the effect on out-of-domain pre-trained models (e.g. ImageNet [60], CLIP [59]), and in-domain pre-trained models (e.g. SatlasNet [5], SeCo [48]). We find that out-of-domain pre-trained models have better downstream performance by applying scale and shift fine-tuning [21, 40], or by applying DoRa [49]. By fine-tuning specific layers, the models retain some of the original representations learned from the pre-training (i.e. ImageNet or CLIP) so that performance does not deteriorate compared to the base models. This has a significant impact on large models such as ViT, since fine-tuning all weights alters many parameters and risks shifting them in suboptimal directions. On the other hand, while applying PEFT for in-domain pre-trained models SatlasNet [5] and SeCo [48] improves performance compared to the base model, we see better performance when directly fine-tuning all the layers. This may be because the model has already undergone pre-training on satellite images, making additional pre-training on similar data from WildSAT result in a non-disruptive shift.